OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 2x + y = -3 5x + 3y = -10

OpenStudy (stefrheart):

ok so first you need to make the x equal to each other

OpenStudy (stefrheart):

try multiplying the top equation by 5

OpenStudy (stefrheart):

do you understand?

OpenStudy (cwrw238):

you could multiply the top equation by 5 and the second equation by -2 to make the terms in x = 10x and -10x then adding would eliminate x and you can then find y

OpenStudy (cwrw238):

remember when you multiply an equation you have to multiply every term

OpenStudy (anonymous):

????

OpenStudy (stefrheart):

what you need o do is multiply the top and bottom

OpenStudy (stefrheart):

on the top you need to multiply it by 5 and the bottom by 2 after that tell me

jhonyy9 (jhonyy9):

you dont understand it ?

jhonyy9 (jhonyy9):

2x+y= -3 *5 5x+3y= -10 *(-2) ---------------------- 10x+5y= -15 -10x -6y = 20 -------------------- 0 - y = 5 y = - 5

jhonyy9 (jhonyy9):

hope you understand it now